OpenStudy (anonymous):

A 2.10-L vessel contains 4.65 g of a gas at 1.00 atm and 27.0°C. (a) Calculate the density of the gas in grams per liter. (b) What is the molar mass of the gas?

OpenStudy (anonymous):

moles gas = pV/RT = 1.00 x 2.10 / 0.08206 x 300=0.0853 molar mass = 4.65 g/ 0.0853 Mol =54.5 g/Mol d = molar mass x p / RT = 54.5 x 1.00 / 0.08206 x 300=2.21 g/L